Based on the Nielsen ratings, the local CBS affiliate claims its 11 p.m. newscast reaches 41% of the viewing audience in the area. In a survey of 100 viewers, 36% indicated that they watch the late evening news on this local CBS station. What is the p-value?
Answer #1

Test is two tailed.

Test statistic formula for this test is given as below:

Z = (p̂ - p)/sqrt(pq/n)

Where, p̂ = Sample proportion, p is population proportion, q = 1 - p, and n is sample size

n = sample size = 100

p̂ = x/n = 0.36

p = 0.41

q = 1 - p = 0.59

Z = (p̂ - p)/sqrt(pq/n)

Z = (0.36 - 0.41)/sqrt(0.41*0.59/100)

Z = -1.0166

Test statistic = -1.0166

P-value = 0.3093

(by using z-table)